# Matrix Is All You Need

### v1

@@ 4.1.1 Data2Matrix
Raw inputs from disparate domains are cast into a consistent matrix format. Grayscale images become two-dimensional matrices whose entries correspond to pixel intensities $g_{ij}$. RGB images are decomposed into three channel-specific matrices. One-dimensional time series form column vectors concatenated horizontally. Multivariate sequences place each observation vector as a row then transpose into a column-wise matrix ($s_{t}=(s_{t1},s_{t2},...,s_{tn})$). Textual sequences begin by embedding tokens into fixed-length vectors; these embeddings populate the rows of a matrix, which is subsequently transposed to align with the weight-matrix dimensions (Token $t_{pos}=(t_{pos,1},t_{pos,2},...,t_{pos,n})$).
$Graph=\begin{pmatrix}g_{11}&\cdots&g_{1n}\\ \vdots&\ddots&\vdots\\ g_{m1}&\cdots&g_{mn}\end{pmatrix}\quad Series=\begin{pmatrix}s_{11}&\cdots&s_{1n}\\ \vdots&\ddots&\vdots\\ s_{n1}&\cdots&s_{nn}\end{pmatrix}\quad Text=\begin{pmatrix}t_{11}&\cdots&t_{1n}\\ \vdots&\ddots&\vdots\\ t_{n1}&\cdots&t_{nn}\end{pmatrix}$
@@ 4.1.2 Matrix Structure
$Graph=\begin{pmatrix}g_{11}&\cdots&g_{1n}\\ \vdots&\ddots&\vdots\\ g_{m1}&\cdots&g_{mn}\end{pmatrix}\quad Graph_{new}=\begin{pmatrix}g_{11}\\ g_{12}\\ g_{13}\\ \vdots\\ g_{mn}\end{pmatrix}$

### v2

@@ 4.1.1 Data2Matrix
Raw inputs from disparate domains are cast into a consistent matrix format. Grayscale images become two-dimensional matrices whose entries correspond to pixel intensities $I_{ij}$. RGB images are decomposed into three channel-specific matrices. One-dimensional time series form column vectors concatenated horizontally. Multivariate sequences place each observation vector as a row then transpose into a column-wise matrix ($s_{t}=(s_{t1},s_{t2},...,s_{tn})$). Textual sequences begin by embedding tokens into fixed-length vectors; these embeddings populate the rows of a matrix, which is subsequently transposed to align with the weight-matrix dimensions (Token $t_{pos}=(t_{pos,1},t_{pos,2},...,t_{pos,n})$).
$Image=\begin{pmatrix}I_{11}&\cdots&I_{1n}\\ \vdots&\ddots&\vdots\\ I_{m1}&\cdots&I_{mn}\end{pmatrix}\quad Series=\begin{pmatrix}s_{11}&\cdots&s_{1n}\\ \vdots&\ddots&\vdots\\ s_{n1}&\cdots&s_{nn}\end{pmatrix}\quad Text=\begin{pmatrix}t_{11}&\cdots&t_{1n}\\ \vdots&\ddots&\vdots\\ t_{n1}&\cdots&t_{nn}\end{pmatrix}$
@@ 4.1.2 Matrix Structure
$Image=\begin{pmatrix}I_{11}&\cdots&I_{1n}\\ \vdots&\ddots&\vdots\\ I_{m1}&\cdots&I_{mn}\end{pmatrix}\quad Image_{new}=\begin{pmatrix}g_{11}\\ g_{12}\\ g_{13}\\ \vdots\\ g_{mn}\end{pmatrix}$